# Bridging Quantum Computing and Nuclear Structure: Atomic Nuclei on a Trapped-Ion Quantum Computer

### v3

@@ I Introduction
The first part, the choice of the ansatz, is particularly crucial. This is because the success of most of quantum algorithms for quantum many-body problems largely depends on the quality of the prepared quantum state, i.e., the overlap between the prepared state and the target state. In spin systems or chemical systems, various ansatze have been proposed, such as UCC-family, and ones based on matrix product states or tensor network states more generally [6, 7, 8]. Compared to other areas tackling quantum many-body problems, nuclear physics studies have seen relatively limited applications of quantum computing, despite several pioneering works [9, 10, 11], and recent reviews [12, 13]. To date, only a few ansatze have been explored for nuclear systems, almost exclusively based on the unitary coupled cluster (UCC) family [14, 15, 11] with some variations such as Adapt-VQE [16]. A seminal work investigated the applicability of the UCC + Adapt-VQE to the nuclear shell model [11], where the number of CNOT gates required to achieve a given accuracy for shell-model calculations of oxygen and calcium isotopes was estimated.
@@
In this work, we introduce a hard-core-boson (HCB) mapping of the nuclear shell model combined with the pair-unitary coupled-cluster doubles (pUCCD) ansatz, which efficiently captures strong pairing correlations between nucleons of the same species. While pUCCD ansatze have been shown to be effective for chemical systems [21, 22, 23], this study reports, to our knowledge, the first demonstration of pUCCD applied to nuclear many-body systems on a quantum device. Implemented on the high-fidelity trapped-ion processor RIKEN-Quantinuum Reimei, and augmented by symmetry-aware ansatze and post-selection techniques, our approach yields ground-state energies for oxygen, calcium, and nickel isotopes that agree with noise-free statevector simulations within sub-percent relative error. These results establish a new experimental benchmark for quantum simulations of strongly correlated nuclear systems and demonstrate the near-term potential of trapped-ion platforms for nuclear-structure calculations.
@@ II.1 Target systems and Hamiltonian mapping
These SPEs and TBMEs can be either derived from the realistic nucleon-nucleon interactions or from the phenomenological adjustments to reproduce experimental data for binding energies, excitation energies, and so on. The target nuclei, model space and adopted interactions in this work are summarized in Table 1. The methods in this work can be applied to any other interactions including microscopically derived ones as far as one considers similar systems with same-species nucleons under the methods described below. Diagonalization with these interactions can be reproduced using publicly available codes such as KSHELL [24, *KSHELL2] or NuclearToolkit.jl [26, *Repo_NuclearToolkit.jl].
@@
The hard-core-boson (HCB) representation provides a natural way to introduce paired degrees of freedom by folding each time-reversed single-particle pair $(i,\bar{i})$ into a single boson-like site that can be occupied or empty but forbids multiple occupancy. Figure 1 illustrates this mapping for 22O on top of the 16O inert core: the valence single-particle states are grouped into pair orbitals, which reduces the number of active degrees of freedom (and hence qubits). The single-particle states in the original fermionic representation are specified by the quantum numbers $\{n,l,j,j_{z},t_{z}\}$, leading to 12 single-particle states for both protons and neutrons in the $sd$ shell. In the HCB representation, these 12 single-particle states form 6 pair orbitals for neutrons, each represented by a qubit. Focusing on even-number neutron systems and terms involving time-reversed pairs, the shell-model Hamiltonian can be expressed as
@@
The paired-neutron operators satisfy bosonic commutation relations, as detailed in e.g. [33, 21, 20], but multiple occupancy of the same pair is forbidden, leading to the term hard-core boson. A key advantage of the HCB representation—beyond the qubit reduction—is the elimination of non-locality. From Jordan-Wigner mappings, Pauli-$Z$ terms appear to maintain the anti-commutation relations of the fermionic operators. In some literature, these Pauli-$Z$ terms are omitted on purpose, which is referred to as qubit excitation in Refs. [36, 37], as a hardware-efficient mapping. However, in the HCB representation, those Pauli-$Z$ terms naturally cancel out. Hence, one can map the Hamiltonian to the Pauli operators into simple forms as
@@
In this work, we employ the pair unitary coupled cluster doubles (pUCCD) ansatz for even-even multi-neutron systems within the HCB representation. The pUCCD ansatz is a quantum version of the pair coupled cluster doubles method. A seminal work on the pUCCD ansatz in the context of quantum computing is found in Ref. [21], where strategies for measurement and post-selection were also proposed. In chemistry, pUCCD has been successfully applied to systems of paired electrons with spin up and spin down [21, 22, 23]. For nuclear systems, the pUCCD ansatz has not been explored previously (to our knowledge), and our work represents the first application of this approach to nuclear structure calculations.
@@
In what follows, we use a single deterministic realization for both pUCCD(GS) and pUCCD(G) unless otherwise noted. The qubits are aligned by descending order of $j$, total angular momentum of the corresponding single-particle states. This is not the optimal lowest-filling configuration; however, we verified that initializing qubits strictly in ascending SPE order is not necessarily superior. The present ordering is therefore a simple and reasonable choice among equivalent alternatives. One can further confirm that this choice does not give an outlier result by comparing with multiple random realizations, as will be shown later in Sec. III.
@@
Table 2: Dimension and number of parameters in ansatze. Comparison of FCI dimension and the number of parameters in the circuits for UCCD($\nu=0$[15] and pUCCD ansatz. The UCCD($\nu=0$) is unitary coupled cluster doubles ansatz where the configurations are restricted to zero-seniority configurations. The rightmost column means the pair unitary coupled cluster doubles ansatz with a single Trotter step.
Nucleus FCI (M=0) UCCD($\nu=0$) pUCCD
18O, 26O 14 5 5
@@ II.4 Implementation of Givens rotation gates
Quantum circuits submitted to the hardware are transformed to native gate sets of the adopted hardware. On the RIKEN-Quantinuum’s Reimei device, two-qubit native gates are ZZPhase gates, $\exp(-i\pi\alpha/2(Z_{p}\otimes Z_{q}))$. Here we follow the notation of pytket [44, *pytket-quantinuum], i.e., angle parameter is given in the unit of $\pi/2$. In practice, the compiled circuit uses exactly twice as many ZZPhase gates as there are Givens rotations.
@@ III.1 Ideal simulation results of the pUCCD ansatz
Table 4: Summary of pUCCD results on the RIKEN-Quantinuum Reimei device. FCI and ideal pUCCD are classical results. Reimei columns report hardware energies using Hadamard or basis-rotation measurements; errors are relative to ideal pUCCD. Valid ratio is the fraction of bitstrings passing particle-number post-selection.
@@
Nuclide Energy (MeV) Energy (MeV) Energy (MeV) Rel. Error (%) Energy (MeV) Rel. Error (%) Valid ratio (%)
18O 11.932-11.932 11.932-11.932 11.802-11.802 $1.09$ 11.893-11.893 $0.33$ $98$
20O 23.632-23.632 22.939-22.939 22.383-22.383 $2.42$ 22.854-22.854 $0.37$ $98$
22O 34.498-34.498 33.924-33.924 33.560-33.560 $1.07$ 33.829-33.829 $0.28$ $97$
24O 41.225-41.225 41.021-41.021 40.985-40.985 $0.09$ 41.027-41.027 $0.02$ $97$
26O 40.869-40.869 40.869-40.869 40.613-40.613 $0.63$ 40.940-40.940 $0.18$ $96$
42Ca 19.734-19.734 19.734-19.734 19.727-19.727 $0.04$ 19.775-19.775 $0.21$ $97$
44Ca 38.675-38.675 38.462-38.462 38.540-38.540 $0.20$ 38.424-38.424 $0.10$ $97$
46Ca 56.667-56.667 56.440-56.440 56.414-56.414 $0.05$ 56.565-56.565 $0.22$ $96$
48Ca 73.662-73.662 73.556-73.556 73.538-73.538 $0.02$ 73.489-73.489 $0.09$ $95$
50Ca 85.055-85.055 84.921-84.921 85.135-85.135 $0.25$ 85.007-85.007 $0.10$ $94$
52Ca 95.360-95.360 95.190-95.190 95.038-95.038 $0.16$ 95.149-95.149 $0.04$ $94$
54Ca 102.632-102.632 102.557-102.557 102.622-102.622 $0.06$ 102.648-102.648 $0.09$ $94$
56Ca 104.589-104.589 104.462-104.462 104.620-104.620 $0.15$ 104.325-104.325 $0.13$ $94$
58Ca 106.666-106.666 106.666-106.666 106.885-106.885 $0.20$ 106.668-106.668 <0.01<0.01 $94$
58Ni 21.447-21.447 21.447-21.447 21.207-21.207 $1.12$ 21.291-21.291 $0.73$ $93$
60Ni 41.276-41.276 40.961-40.961 40.752-40.752 $0.51$ 40.992-40.992 $0.07$ $94$
62Ni 59.602-59.602 59.013-59.013 58.703-58.703 $0.52$ 59.015-59.015 <0.01<0.01 $93$
64Ni 76.444-76.444 75.687-75.687 76.027-76.027 $0.45$ 75.765-75.765 $0.10$ $92$
66Ni 91.941-91.941 91.059-91.059 90.552-90.552 $0.56$ 90.695-90.695 $0.40$ $93$
68Ni 106.140-106.140 105.237-105.237 105.169-105.169 $0.06$ 105.288-105.288 $0.05$ $92$
70Ni 119.034-119.034 118.215-118.215 118.189-118.189 $0.02$ 118.158-118.158 $0.05$ $91$
72Ni 130.930-130.930 130.379-130.379 130.511-130.511 $0.10$ 130.397-130.397 $0.01$ $95$
74Ni 141.932-141.932 141.694-141.694 141.565-141.565 $0.09$ 141.715-141.715 $0.02$ $93$
76Ni 152.103-152.103 152.103-152.103 151.986-151.986 $0.08$ 152.054-152.054 $0.03$ $94$
Fig. 3 shows ground-state energies for the target isotopes, obtained with pUCCD(G) and pUCCD(GS) on noise-free statevector simulators compared to FCI and DOCI results. The upper panels (a)–(c) display the ground-state energies. At this scale, all curves are nearly indistinguishable. The lower panels (d)–(f) show the differences between pUCCD and DOCI results relative to FCI in logarithmic scale. Results for two-neutron (or two-neutron-hole) systems are omitted in the lower panels, since both DOCI and pUCCD exactly reproduce FCI in those cases.
Most importantly, pUCCD reproduces DOCI energies with an error of mostly less than $1\%$, while using far fewer parameters, demonstrating its compact yet accurate description. For reference, DOCI itself reproduces FCI ground-state energies within $0.1-1\%$, with deviations smaller than the typical uncertainties (a few percent) of modern nuclear many-body methods [47]. The residual differences depend on the target system: Ca isotopes in the $pf$ shell are well described within the zero-seniority space, whereas oxygen isotopes show larger contributions from higher-seniority configurations leading to slightly larger deviations from FCI. Nickel isotopes in the $jj45$ shell exhibit larger discrepancies between DOCI and pUCCD compared to oxygen and calcium isotopes. This arises from the factors intrinsic to the $jj45$ shell. The valence space of the $jj45$ shell consists of four $jj$-coupling orbitals ($1p_{1/2},1p_{3/2},0f_{5/2},0g_{9/2}$), and the configuration mixing among these orbitals is pronounced in Ni isotopes even for ground states since their effective single particle energies are close to each other [30]. The single-Trotter pUCCD ansatz, while efficient for capturing pairing correlations, becomes less expressive in the presence of such strong configuration mixing.
@@ IV Summary and outlook
Although the pUCCD parameters were optimized using classical simulators, this proof-of-principle study highlights how physically motivated, compact ansatze can deliver accurate results with modest qubit numbers and circuit depths. Such strategies may guide the discovery of efficient variational forms for larger systems where classical pre-optimization is infeasible.
Looking ahead, incorporating proton-neutron interactions will require ansatze capable of describing more general non-local correlations beyond pairing. Nonetheless, the present results demonstrate the viability of pUCCD for multi-neutron (or multi-proton) systems, and point toward applications to heavier systems such as neutron drops or neutron-rich isotopes. As an illustration of possible extensions, entropy-driven entanglement forging [48] has recently been proposed to localize correlations in Adapt-VQE circuits; similar ideas could be combined with pUCCD, for instance by monitoring generalized seniority. In addition, hybrid qubit encoding schemes [49] could be extended to nuclear systems.
@@
The data presented in this manuscript, including circuits and encoded Hamiltonians, are available from the GitHub repository and the Zenodo archive [55]. For classical calculations, the exact diagonalization results can be reproduced by using the NuclearToolkit.jl package [26, *Repo_NuclearToolkit.jl] with the interaction files available through KSHELL [24, *KSHELL2]. We utilized the Qiskit [56], PennyLane [39], and pytket [44, *pytket-quantinuum] for quantum computations. Experiments on the Reimei device were carried out using the Quantinuum’s Nexus [46] platform.

### v4

@@ I Introduction
The first part, the choice of the ansatz, is particularly crucial. This is because the success of most of quantum algorithms for quantum many-body problems largely depends on the quality of the prepared quantum state, i.e., the overlap between the prepared state and the target state. In spin systems or chemical systems, various ansatze have been proposed, such as UCC-family, and ones based on matrix product states or tensor network states more generally [6, 7, 8]. Compared to other areas tackling quantum many-body problems, nuclear physics studies have seen relatively limited applications of quantum computing, despite several pioneering works [9, 10, 11], and recent reviews [12, 13]. To date, only a few ansatze have been explored for nuclear systems, almost exclusively based on the unitary coupled cluster (UCC) family [14, 15, 11] with some variations such as Adapt-VQE [16]. A seminal work investigated the applicability of the UCC + Adapt-VQE to the nuclear shell model [11], where the number of CNOT gates required to achieve a given accuracy for shell-model calculations of oxygen and calcium isotopes was estimated. Taking a look around existing literature, pioneering works have focused intensively on pairing Hamiltonians [10, 17, 18, 19, 20, 21, 22]. Building on these previous works on pairing Hamiltonians or pairing-like ansatze, this work aims to explore more expressive and scalable ansatze suitable for simplified representations of nuclear shell-model Hamiltonians, develop symmetry-aware post-selection techniques, and evaluate their performance on quantum hardware across a range of nuclear systems.
@@
In this work, we introduce a hard-core-boson (HCB) mapping of the nuclear shell model combined with the pair-unitary coupled-cluster doubles (pUCCD) ansatz, which efficiently captures strong pairing correlations between nucleons of the same species. While pUCCD ansatze have been shown to be effective for chemical systems [27, 28, 29], applications of such pair-based ansatze to nuclear systems have been limited a few pioneering studies: [19, 22, 30]. This work extends these previous studies to shell-model Hamiltonians, and reports the first hardware implementation of HCB and pUCCD for nuclear systems. Implemented on the high-fidelity trapped-ion processor RIKEN-Quantinuum Reimei, and augmented by symmetry-aware ansatze and post-selection techniques, our approach yields ground-state energies for oxygen, calcium, and nickel isotopes that agree with noise-free statevector simulations within sub-percent relative error. These results establish a new experimental benchmark for quantum simulations of strongly correlated nuclear systems and demonstrate the near-term potential of trapped-ion platforms for nuclear-structure calculations.
@@ II.1 Target systems and Hamiltonian mapping
These SPEs and TBMEs can be either derived from the realistic nucleon-nucleon interactions or from the phenomenological adjustments to reproduce experimental data for binding energies, excitation energies, and so on. The target nuclei, model space and adopted interactions in this work are summarized in Table 1. The methods in this work can be applied to any other interactions including microscopically derived ones as far as one considers similar systems with same-species nucleons under the methods described below. Diagonalization with these interactions can be reproduced using publicly available codes such as KSHELL [31, *KSHELL2] or NuclearToolkit.jl [33, *Repo_NuclearToolkit.jl].
@@
The hard-core-boson (HCB) representation provides a natural way to introduce paired degrees of freedom by folding each time-reversed single-particle pair $(i,\bar{i})$ into a single boson-like site that can be occupied or empty but forbids multiple occupancy. Similar mappings have been employed in various contexts within quantum chemistry and nuclear physics [42, 43, 27, 17, 19, 26, 22, 30]. Depending on whether one emphasizes the underlying degrees of freedom or the commutation relations of the operators, this construction is referred to by various names—geminal, hard-core-boson (HCB), antisymmetrized geminal power (AGP), seniority-zero, quasi-particle representations, etc.—but the underlying idea is similar.
Here, we summarize the derivation of the HCB representation for nuclear shell-model Hamiltonians. Figure 1 illustrates this mapping for 22O on top of the 16O inert core: the valence single-particle states are grouped into pair orbitals, which reduces the number of active degrees of freedom (and hence qubits). The single-particle states in the original fermionic representation are specified by the quantum numbers $\{n,l,j,j_{z},t_{z}\}$, leading to 12 single-particle states for both protons and neutrons in the $sd$ shell. In the HCB representation, these 12 single-particle states form 6 pair orbitals for neutrons, each represented by a qubit. Focusing on even-number neutron systems and terms involving time-reversed pairs, the shell-model Hamiltonian can be expressed as
@@
The paired-neutron operators satisfy bosonic commutation relations, as detailed in e.g. [40, 27, 26], but multiple occupancy of the same pair is forbidden, leading to the term hard-core boson. A key advantage of the HCB representation—beyond the qubit reduction—is the elimination of long strings of Pauli-$Z$ operators. From Jordan-Wigner mappings, Pauli-$Z$ terms appear to maintain the anti-commutation relations of the fermionic operators. In some literature, these Pauli-$Z$ terms are omitted on purpose, which is referred to as qubit excitation in Refs. [45, 46], as a hardware-efficient mapping. However, in the HCB representation, those Pauli-$Z$ terms naturally cancel out. Hence, one can map the Hamiltonian to the Pauli operators into simple forms as
@@
In this work, we employ the pair unitary coupled cluster doubles (pUCCD) ansatz within the HCB representation, which is particularly well-suited for systems with an even number of valence neutrons. The pUCCD ansatz is a quantum version of the pair coupled cluster doubles method. A seminal work on the pUCCD ansatz in the context of quantum computing is found in Ref. [27], where strategies for measurement and post-selection were also proposed. In chemistry, pUCCD has been successfully applied to systems of paired electrons with spin up and spin down [27, 28, 29]. For nuclear systems, the pUCCD ansatz has been considered in a few theoretical studies [19], but to our knowledge, no experimental implementation has been reported so far.
@@
In what follows, we use a single deterministic realization for both pUCCD(GS) and pUCCD(G) unless otherwise noted. The qubits are aligned by ascending order of $j$, total angular momentum of the corresponding single-particle states. This is not the optimal lowest-filling configuration; however, we verified that initializing qubits strictly in ascending SPE order is not necessarily superior. The present ordering is therefore a simple and reasonable choice among equivalent alternatives. One can further confirm that this choice does not give an outlier result by comparing with multiple random realizations, as will be shown later in Sec. III.
@@
Table 2: Comparison of Hilbert-space dimension and ansatz parameter counts. For each nucleus we list: the number of qubits $N_{q}$ in both Fermionic and HCB representations, the full configuration-interaction (FCI) dimension in the $M=0$ sector, the number of variational parameters required by UCCD($\nu=0$) (unitary coupled-cluster doubles restricted to zero-seniority/paired configurations), and the number of parameters for the pUCCD ansatz implemented with a single Trotter step.
$N_{q}$ Dimensions
Nucleus Fermionic/HCB FCI UCCD($\nu=0$) pUCCD
18O, 26O 12/6 14 5 5
@@ II.4 Implementation of Givens rotation gates
Quantum circuits submitted to the hardware are transformed to native gate sets of the adopted hardware. On the RIKEN-Quantinuum’s Reimei device, two-qubit native gates are ZZPhase gates, $\exp(-i\pi\alpha/2(Z_{p}\otimes Z_{q}))$. Here we follow the notation of pytket [53, *pytket-quantinuum], i.e., angle parameter is given in the unit of $\pi/2$. In practice, the compiled circuit uses exactly twice as many ZZPhase gates as there are Givens rotations.
@@ III.1 Ideal simulation results of the pUCCD ansatz
Fig. 3 shows ground-state energies for the target isotopes, obtained with pUCCD(G) and pUCCD(GS) on noise-free statevector simulators compared to FCI and DOCI results. The upper panels (a)–(c) display the ground-state energies. At this scale, all curves are nearly indistinguishable. The lower panels (d)–(f) show the differences between pUCCD and DOCI results relative to FCI in logarithmic scale. Results for two-neutron (or two-neutron-hole) systems are omitted in the lower panels where both DOCI and pUCCD exactly reproduce FCI in these cases. This is because the HCB representation with paired operators can e xactly represent the ground states of single pair systems without approximation [26].
Most importantly, pUCCD reproduces DOCI energies with an error of mostly less than $1\%$, while using far fewer parameters, demonstrating its compact yet accurate description. For reference, DOCI itself reproduces FCI ground-state energies within $0.1-1\%$, with deviations smaller than the typical uncertainties (a few percent) of modern nuclear many-body methods [56]. The residual differences depend on the target system: Ca isotopes in the $pf$ shell are well described within the zero-seniority space, whereas oxygen isotopes show larger contributions from higher-seniority configurations leading to slightly larger deviations from FCI. Nickel isotopes in the $jj45$ shell exhibit larger discrepancies between DOCI and pUCCD compared to oxygen and calcium isotopes. This arises from the factors intrinsic to the $jj45$ shell. The valence space of the $jj45$ shell consists of four $jj$-coupling orbitals ($1p_{1/2},1p_{3/2},0f_{5/2},0g_{9/2}$), and the configuration mixing among these orbitals is pronounced in Ni isotopes even for ground states since their effective single particle energies are close to each other [37]. The results imply that the dominant pairing correlations in the same-species nucleons are expressed rather separable manner, and that one may need to consider more expressive ansatze, e.g., by increasing the number of Trotter steps or including higher-order excitations, to further improve the accuracy for systems with strong configuration mixing.
[…]
@@
Nuclide Energy Energy Energy std. Rel. Error Energy std. Rel. Error Valid ratio (%)
18O 11.932-11.932 11.932-11.932 11.802-11.802 (0.209) $1.09$ 11.893-11.893 (0.099) $0.33$ $0.98$
20O 23.632-23.632 22.939-22.939 22.383-22.383 (0.234) $2.42$ 22.854-22.854 (0.140) $0.37$ $0.98$
22O 34.498-34.498 33.924-33.924 33.560-33.560 (0.222) $1.07$ 33.829-33.829 (0.152) $0.28$ $0.97$
24O 41.225-41.225 41.021-41.021 40.985-40.985 (0.194) $0.09$ 41.027-41.027 (0.150) $0.02$ $0.97$
26O 40.869-40.869 40.869-40.869 40.613-40.613 (0.176) $0.63$ 40.940-40.940 (0.106) $0.18$ $0.96$
42Ca 19.734-19.734 19.734-19.734 19.727-19.727 (0.126) $0.04$ 19.775-19.775 (0.053) $0.21$ $0.97$
44Ca 38.675-38.675 38.462-38.462 38.540-38.540 (0.162) $0.20$ 38.424-38.424 (0.073) $0.10$ $0.97$
46Ca 56.667-56.667 56.440-56.440 56.414-56.414 (0.143) $0.05$ 56.565-56.565 (0.062) $0.22$ $0.96$
48Ca 73.662-73.662 73.556-73.556 73.538-73.538 (0.131) $0.02$ 73.489-73.489 (0.080) $0.09$ $0.95$
50Ca 85.055-85.055 84.921-84.921 85.135-85.135 (0.151) $0.25$ 85.007-85.007 (0.087) $0.10$ $0.94$
52Ca 95.360-95.360 95.190-95.190 95.038-95.038 (0.146) $0.16$ 95.149-95.149 (0.095) $0.04$ $0.94$
54Ca 102.632-102.632 102.557-102.557 102.622-102.622 (0.127) $0.06$ 102.648-102.648 (0.081) $0.09$ $0.94$
56Ca 104.589-104.589 104.462-104.462 104.620-104.620 (0.143) $0.15$ 104.325-104.325 (0.082) $0.13$ $0.94$
58Ca 106.666-106.666 106.666-106.666 106.885-106.885 (0.132) $0.20$ 106.668-106.668 (0.058) <0.01<0.01 $0.94$
58Ni 21.447-21.447 21.447-21.447 21.207-21.207 (0.149) $1.12$ 21.291-21.291 (0.071) $0.73$ $0.93$
60Ni 41.276-41.276 40.961-40.961 40.752-40.752 (0.197) $0.51$ 40.992-40.992 (0.094) $0.07$ $0.94$
62Ni 59.602-59.602 59.013-59.013 58.703-58.703 (0.228) $0.52$ 59.015-59.015 (0.110) <0.01<0.01 $0.93$
64Ni 76.444-76.444 75.687-75.687 76.027-76.027 (0.261) $0.45$ 75.765-75.765 (0.116) $0.10$ $0.92$
66Ni 91.941-91.941 91.059-91.059 90.552-90.552 (0.245) $0.56$ 90.695-90.695 (0.138) $0.40$ $0.93$
68Ni 106.140-106.140 105.237-105.237 105.169-105.169 (0.263) $0.06$ 105.288-105.288 (0.133) $0.05$ $0.92$
70Ni 119.034-119.034 118.215-118.215 118.189-118.189 (0.253) $0.02$ 118.158-118.158 (0.115) $0.05$ $0.91$
72Ni 130.930-130.930 130.379-130.379 130.511-130.511 (0.223) $0.10$ 130.397-130.397 (0.092) $0.01$ $0.95$
74Ni 141.932-141.932 141.694-141.694 141.565-141.565 (0.197) $0.09$ 141.715-141.715 (0.074) $0.02$ $0.93$
76Ni 152.103-152.103 152.103-152.103 151.986-151.986 (0.133) $0.08$ 152.054-152.054 (0.053) $0.03$ $0.94$
@@ IV Summary and outlook
While the pUCCD parameters in this work were optimized on classical simulators, this proof-of-principle study shows that compact, physically motivated ansatze can achieve high accuracy with modest qubit counts and circuit depths. Such strategies can guide the design of efficient variational forms for larger systems where classical pre-optimization is infeasible. Optimization for classically intractable systems remains an open challenge, and new, hardware-suitable optimization methods are needed. Our previous study [26] demonstrated that a sequential, gradient-free optimization scheme [49] can effectively optimize a similar ansatz for two-valence-neutron systems. Although those systems were still classically tractable, the results suggest that sequential methods may be viable for optimizing or refining ansatze on quantum hardware. For larger problems, hybrid approaches that combine classical pre-optimization in a reduced subspace with subsequent quantum refinement could also be explored.
It remains an open question how best to extend the present framework to include proton–neutron interactions, which are essential for a complete description of nuclei. While several studies have investigated proton–neutron pairing alongside like-particle pairing [19, 22], a practical quantum ansatz must be able to capture more general, nonlocal correlations beyond simple pair excitations. Developing such ansatze and evaluating their resource and measurement requirements will be necessary to treat mixed-species correlations robustly. Nonetheless, our results demonstrate that pUCCD provides a compact and accurate description of pairing correlations in single-species systems (multi-neutron or multi-proton), establishing a foundation for future extensions to more complex nuclear systems, and footsteps toward heavier systems such as neutron drops or neutron-rich isotopes.
As an illustration of possible extensions, entropy-driven entanglement forging [57] has recently been proposed to localize correlations in Adapt-VQE circuits; similar ideas including Ref. [58] could be combined with pUCCD, for instance by monitoring generalized seniority. In addition, hybrid qubit encoding schemes [59] could be extended to nuclear systems.
@@
The data presented in this manuscript, including circuits and encoded Hamiltonians, are available from the GitHub repository and the Zenodo archive [65]. For classical calculations, the exact diagonalization results can be reproduced by using the NuclearToolkit.jl package [33, *Repo_NuclearToolkit.jl] with the interaction files available through KSHELL [31, *KSHELL2]. We utilized the Qiskit [66], PennyLane [48], and pytket [53, *pytket-quantinuum] for quantum computations. Experiments on the Reimei device were carried out using the Quantinuum’s Nexus [55] platform.